OpenStudy (anonymous):

How do you find the derivative of ln(x+6)

OpenStudy (turingtest):

do you know the derivative of lnx ?

OpenStudy (anonymous):

If You Don't* : Derivative of lnx= (1/x)*(derivative of x) example: Find derivative of ln2x d(ln2x)/dx = (1/2x)*d(2x)/dx = (1/2x)*2===>1/x

OpenStudy (anonymous):

Step 1: Find the derivative of ln(x+6) hint: treat it as ln(x) Step 2: then multiply by the derivative of inside of the parenthesis .... this last step shouldn't matter much though

OpenStudy (anonymous):

So would it be 6/x?

OpenStudy (turingtest):

no, do you know the chain rule?

OpenStudy (turingtest):

@ally7343 ^

OpenStudy (anonymous):

i forgot how to do it...

OpenStudy (anonymous):

I got it, thanks anyways